Exercise 2-17 (Part Level Submission) Erin Brushwood sells gourmet chocolate chip cookies. The results of her last month of operations are as follows: Sales revenue $ 49,742 Cost of goods sold (all variable) 25,443 Gross margin 24,299 Selling expenses (20% variable) 8,239 Administrative expenses (60% variable) 11,152 Operating income $ 4,908 Collapse question part (a) Partially correct answer. Your answer is partially correct. Prepare a contribution format income statement for Erin. (Round answers to 0 decimal places, e.g. 5,275.) Entry field with correct answer $Entry field with correct answer 49742 Entry field with correct answer: Entry field with correct answer $ Entry field with correct answer 25443 Entry field with correct answer Entry field with correct answer 1648 Entry field with incorrect answer Entry field with correct answer 6691 Entry field with incorrect answer Entry field with correct answer 33782 Entry field with correct answer Entry field with correct answer 15960 Entry field with correct answer: Entry field with correct answer Entry field with correct answer 6591 Entry field with correct answer Entry field with correct answer 4461 Entry field with correct answer Entry field with correct answer 11052 Entry field with correct answer $Entry field with correct answer 4908 Click if you would like to Show Work for this question: Open Show Work SHOW SOLUTION LINK TO TEXT LINK TO VIDEO Attempts: 1 of 1 used Collapse question part (b) Correct answer. Your answer is correct. If Erin sells her cookies for $1.60 each, how many cookies did she sell during the month? (Round answer to the nearest whole number, e.g. 5,275.) Cookies sold in the month Entry field with correct answer 31089 Click if you would like to Show Work for this question: Open Show Work SHOW SOLUTION LINK TO TEXT LINK TO VIDEO Attempts: 1 of 1 used Collapse question part (c) What is the contribution margin per cookie? (Round unit costs to 2 decimal places, e.g. 52.75.) Contribution margin $ per cookie
Answer #1
a
Sales revenue 49742
Variable expenses:
Cost of goods sold 25443
Selling expense 1648 =8239*20%
Administrative expense 6691 =11152*60%
Total Variable expenses 33782
Contribution margin 15960
Fixed expenses:
Selling expense 6591 =8239*80%
Administrative expense 4461 =11152*40%
Total Fixed expenses 11052
Operating income 4908
b
Cookies sold in the month 31089 =49742/1.60
c
Contribution margin 0.51 per cookie =15960/31089